Find (x̅, y̅): R = {(x, y): 0 ≤ y ≤ √x^2 + 1 , 0 ≤ x ≤ 1} about the x-axis.


1
Expert's answer
2022-05-30T23:25:44-0400

When a curve "y=f(x)" is rotated about the "x" - axis, the centre of mass of the solid generated will lie on the "x" – axis because of the symmetry of the curve.


"\bar{x}=\dfrac{\displaystyle\int_{0}^{1}xy^2dx}{\displaystyle\int_{0}^{1}y^2dx}=\dfrac{\displaystyle\int_{0}^{1}(x^3+x)dx}{\displaystyle\int_{0}^{1}(x^2+1)dx}"

"=\dfrac{[\dfrac{x^4}{4}+\dfrac{x^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}}{[\dfrac{x^3}{3}+x]\begin{matrix}
 1 \\
 0
\end{matrix}}=\dfrac{9}{16}"

"(\dfrac{9}{16}, 0)"